40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4183     Entry time: Fri Jan 21 15:26:15 2011
Author: josephb 
Type: Update 
Category: CDS 
Subject: c1sus broken yesterday and now fixed 

[Joe, Koji]
Yesterday's CDS swap of c1sus and c1iscex left the interfometer in a bad state due to several issues.

The first being a need to actually power down the IO chassis completely (I eventually waited for a green LED to stop glowing and then plugged the power back in) when switching computers.  I also plugged and plugged the interface cable from the IO chassis and computer while powered down.  This let the computer actually see the IO chassis (previously the host interface card was glowing just red, no green lights).

Second, the former c1iscex computer and now new c1sus computer only has 6 CPUs, not 8 like most of the other front ends.  Because it was running 6 models (c1sus, c1mcs, c1rms, c1rfm, c1pem, c1x02) and 1 CPU needed to be reserved for the operating system, 2 models were not actually running (recycling mirrors and PEM).  This meant the recycling mirrors were left swinging uncontrolled.

To fix this I merged the c1rms model with the c1sus model.  The c1sus model now controls BS, ITMX, ITMY, PRM, SRM.  I merged the filter files in the /chans/ directory, and reactivated all the DAQ channels.  The master file for the fb in the /target/fb directory had all references to c1rms removed, and then the fb was restarted via "telnet fb 8088" and then "shutdown".

My final mistake was starting the work late in the day.

So the lesson for Joe is, don't start changes in the afternoon.

Koji has been helping me test the damping and confirm things are really running.  We were having some issues with some of the matrix values.  Unfortunately I had to add them by hand since the previous snapshots no longer work with the models.

ELOG V3.1.3-